# FrugalPrompt: Reducing Contextual Overhead in Large Language Models via Token Attribution

### v2

@@ Abstract
Large language models (LLMs) owe much of their stellar performance to expansive input contexts, yet such verbosity inflates monetary costs, carbon footprint, and inference-time latency. Much of this overhead manifests from the redundant low-utility tokens present in typical prompts, as only a fraction of tokens typically carries the majority of the semantic weight. We address this inefficiency by introducing FrugalPrompt, a novel prompt compression framework for LLMs, which retains only the most semantically significant tokens. Leveraging two state-of-the-art token attribution methods, GlobEnc and DecompX, we assign salience scores to every token in an input sequence, rank them to preserve the top-$k\%$ tokens in their original order, and obtain a sparse frugalized prompt. We evaluate the approach across four NLP tasks: Sentiment Analysis, Commonsense QA, Summarization, and Mathematical Reasoning, using a suite of frontier LLMs. For the first three tasks, a 20% prompt reduction incurs only a marginal loss in task performance, demonstrating that contemporary LLMs can reconstruct elided context from high-salience cues. In contrast, performance on mathematical reasoning deteriorates sharply, reflecting a stronger dependence on complete token continuity. Further analysis with bottom-$k\%$ and random-$k\%$ tokens reveals asymmetric performance patterns that may suggest potential task contamination effects, wherein models may resort to shallow memorized patterns from pretraining exposure for conventional NLP tasks. We posit that our work contributes to a more nuanced understanding of LLM behavior in performance-efficiency trade-offs, and delineate the boundary between tasks tolerant to contextual sparsity and those requiring exhaustive context. Our source code and data are available at: https://github.com/Starscream-11813/Frugal-ICL.
Keywords: Large Language Models, Prompt Compression, Token Attribution, Salience Scores, GlobEnc, DecompX, Inference Efficiency, Task Contamination, Contextual Redundancy.
[…]
FrugalPrompt: Reducing Contextual Overhead in Large Language Models via Token Attribution
Syed Rifat Raiyan1  Md Farhan Ishmam2
Abdullah Al Imran3   Mohammad Ali Moni4∗thanks: * Corresponding author.
1Islamic University of Technology, Bangladesh 2University of Utah, USA
3University of Liverpool, UK 4University of Queensland, Australia
rifatraiyan@iut-dhaka.edu farhan.ishmam@utah.edu
abdalimran@gmail.com m.moni@uq.edu.au
Abstract content
1.  Introduction
As large language models (LLMs) continue to reshape the landscape of NLP, they are increasingly applied to complex and ingenious tasks Yang et al. (2024); Schmidgall et al. (2025). With the scaling of LLMs into super-large trillion-parameter models Fedus et al. (2022); Hudson et al. (2023), and multi-agent architectures requiring hundreds of API calls Islam et al. (2024); Jang et al. (2024), the computational footprint has increased drastically Wu et al. (2025), resulting in high inference latency Wan et al. (2023) and energy consumption Stojkovic et al. (2024).
To counter the consequences of the scaling law for LLMs Kaplan et al. (2020), several methods were introduced to improve model efficiency Wan et al. (2023); Zhou et al. (2024). Prompt compression is one of the subfields where the input prompting tokens are reduced while retaining considerable performance Zhang et al. (2025). Compression is either achieved on the hard prompts (natural language input text) Sennrich et al. (2016) or the soft prompt (trainable embedding-size vectors)Zhao et al. (2023a), with the former being more popular due to its interpretability and compatibility with closed-source models.
Current hard prompt compression methods fall short in three key aspects. First, some methods rely on summarizing the input prompt Chuang et al. (2024), which can inadvertently discard semantically critical tokens. Second, these methods tend to retain low-information tokens, such as articles and conjunctions Jiang et al. (2023). Third, all methods either rely on an auxiliary smaller LLM with billions of parameters or a training paradigm. For instance, Chuang et al. (2024) use a Vicuna 7B model Chiang et al. (2023), which introduces a significant parameter overhead.
@@ 1.  Introduction
We address these challenges by introducing a prompt compression strategy that uses saliency scores generated from state-of-the-art pre-trained token attribution methods Modarressi et al. (2022, 2023). These saliency scores can capture the semantically rich token while filtering the redundant ones. This is achieved without any training paradigm, with minimal parameter overhead of a single 110M BERT model Devlin et al. (2019), substantially lower than that of existing methods.
@@
Our criteria for filtering are grounded in two linguistic principles underlying textual redundancy. First, there is an intrinsic redundancy to any form of text produced by humans Wit and Gillette (1999). For instance, in the text, “What are your plans for the weekend, by the way?”, the phrase “by the way” contributes no additional meaning to an LLM. Second, the English language has low-importance tokens that add little to no semantic weight, generally in the form of articles and auxiliary verbs. For instance, most of the meaning is conveyed if the statement “I have received a parcel.”, is simplified to “I received parcel.", with the latter mimicking simplified English, often used by non-native speakers Shubert et al. (1995). Our filtering methods eliminate both forms of redundancy.
Our key contributions are as follows:
@@
  • We introduce a novel, training-free prompt compression strategy for LLMs that controllably filters low-importance tokens based on saliency scores from pre-trained encoders.
  • Experiments on four tasks reveal that a 20% prompt reduction retains the performance across most models and tasks with minimal parameter overhead.
@@
  • We reveal key behavioral insights into the performance scaling of models with the inference cost and possible task contamination in common benchmarks.
2.  Related Work
2.1.  LLM Efficiency
@@
The current literature on LLM efficiency Wan et al. (2023) stems from the literature on efficient transformers Tay et al. (2022), but for the autoregressive setting of LLMs. Inference-time efficiency can be achieved via model compression methods, such as quantization Dettmers et al. (2023b), parameter pruning Ma et al. (2023), low-rank decomposition Li et al. (2023a), and knowledge distillation Timiryasov and Tastet (2023). Training efficiency can be achieved during pre-training Liu et al. (2023); Yao et al. (2024) and fine-tuning, either through parameter-efficient Hu et al. (2022) or memory-efficient Dettmers et al. (2023a) approaches. Finally, architectural changes also improve efficiency, e.g., efficient attention Child et al. (2019); Ainslie et al. (2023) and state-space models Gu and Dao (2023). Distinct from these methods, our work adopts a data-centric approach where the input prompt is optimized instead of changing the underlying weights or architecture of the model Chang et al. (2024).
$2.$  Prompt Compression
Prompt compression aims to minimize input redundancy while retaining task-relevant information. Trainable soft-prompts achieve these through learned embeddings Zhao et al. (2023b), while sacrificing the interpretability of the prompt text. In contrast, compressing hard prompts preserves interpretability via filtering important tokens Li et al. (2023b); Jiang et al. (2023) or paraphrasing the prompt text into a reduced version Chuang et al. (2024). Context-heavy tasks, e.g., conversation systems, benefit from prompt compression but have focused on sentential redundancy or summarization Santra et al. (2023). Our work instead aims at token-level reduction, offering higher granularity without losing semantic content.
While our work primarily filters the top $k\%$ tokens based on the attribution scores, a key distinction is that we use smaller BERT-class text encoders to gain attribution scores. Previous works lack such efficiency as they rely on smaller auto-regressive models, e.g. Jiang et al. (2023) uses GPT-2 Radford et al. (2019) and Chuang et al. (2024) uses Vicuna 7B Chiang et al. (2023). While the compressor model has fewer parameters than the interference model, its overhead is significantly higher than that of the discriminative token attribution methods.
2.3.  Token Attribution
@@
The self-attention weights in transformers have been primarily explored as a proxy for token importance Htut et al. (2019); Clark (2019), yet they fail to align with the true attribution of tokens Jain and Wallace (2019). While including vector norms can partially mitigate this issue Kobayashi et al. (2020), it overlooks the intricacies of the non-attention components in a transformer. Instead, we rely on subsequent token attribution methods: GlobEnc Modarressi et al. (2022), which quantifies global token attribution by aggregating layer-wise analyses across the entire encoder using attention rollout and vector norms, and DecompX Modarressi et al. (2023), which propagates locally decomposed token representations globally to incorporate all model components, including non-linear feed-forward networks. These methods were selected for their superior ability to capture comprehensive semantic contributions beyond raw attention weights, enabling accurate identification of salient tokens with minimal computational overhead.
@@
3.  Preliminaries: Token Attribution
3.1.  GlobEnc
[…]
$\hat{\mathbf{A}}_{l}$ $=0.5\bar{\mathbf{A}}_{l}+0.5\mathbf{I}$ (4)
3.2.  DecompX
where the normalized contribution is given by:
$\mathbf{z}_{\text{FFN}}^{\ell}$ $=\text{FFN}(\tilde{\mathbf{z}}_{i}^{\ell})$ (7)
$y_{c}=\sum_{k=1}^{N}y_{c\Leftarrow k}$ (10)
4.  Methodology
Figure 1 portrays an overview of our proposed pipeline. We obtain the saliency scores using the token attribution methods discussed in Sec. 3 and employ frugalization via token filtration to generate a condensed yet semantically representative version of an input sample. This technique identifies and retains only the most salient tokens, thereby reducing input dimensionality while preserving the core meaning for the given task.
4.1.  Ranking via Saliency Scores
@@ 4.2.  Top-$k$ Filtering
We reindex $S_{k}$ in monotone increasing order to produce ${S}_{k\uparrow}$. The frugalized sequence $\mathbf{F}_{k}$ is produced by restoring the selected tokens to their original order, i.e.,
@@
Model
Token
Attribu-
tion
$k\%$ Tasks
@@
Figure 2: The performance difference between the reduced tokens and baseline performance using 100% tokens of metrics across three tasks: text classification (CLS), question-answering (QA), and reasoning (RSN).
Figure 3: The performance difference between the reduced tokens and baseline performance using 100% tokens across the summarization metrics.
5.  Experiment
5.1.  Tasks
@@
5.2.  Dataset
For sentiment analysis, we use the IMDb movie review dataset Maas et al. (2011), a binary sentiment analysis corpus comprising movie reviews labeled positive or negative. We use the Argilla News dataset Ahmed et al. (2017) to evaluate abstractive summarization on news articles. For commonsense QA, we use the CosmosQA dataset Huang et al. (2019), which primarily focuses on commonsensical and causal reasoning within short narrative passages. The questions stipulated in CosmosQA are easy for humans to solve, but usually challenging for LLMs. We use the GSM8k dataset Cobbe et al. (2021) for mathematical reasoning, which contains grade-school math word problems to assess basic arithmetic reasoning and multi-step logical deduction.
5.3.  Models
We perform inference using a set of contemporary frontier models that vary in architecture and scale. This is primarily an effort to examine the generalizability of our method. The models are: LLaMA-3 8B, LLaMA-3 70B Meta (2024), GPT-3.5 Brown et al. (2020), Gemini 2.0 Flash Thinking Google DeepMind (2025), and o3-mini OpenAI (2025). The closed-source models are accessed via their respective APIs, and the open-source models are accessed via OpenRouter [^1] https://openrouter.ai/, using standardized prompts to minimize formatting-induced variability.
5.4.  Experimental Setup
We set the temperature to $1.0$ for all inference runs to sample from the model’s unaltered probability distribution, ensuring a standard and reproducible generation process. For the frugalized setup, we experiment by filtering the top $k\in\{50,60,80\}\%$ of the input tokens. We also include the baseline prompt using $100\%$ of the tokens to set a proper frame of reference for the performance delta for all the models. This setup facilitates us to investigate whether low-attribution filler tokens can be excluded with minimal degradation in task performance, which, by implication, will also reduce API costs and minimize carbon footprint.
$5.$  Evaluation Metrics

### v3

@@ Abstract
Human communication heavily relies on laconism and inferential pragmatics, allowing listeners to successfully reconstruct rich meaning from sparse, telegraphic speech. In contrast, large language models (LLMs) owe much of their stellar performance to expansive input contexts, yet such verbosity inflates monetary costs, carbon footprint, and inference-time latency. This overhead manifests from the redundant low-utility tokens present in typical prompts, as only a fraction of tokens typically carries the majority of the semantic weight. Inspired by the aforementioned cognitive psycholinguistic processes, we address this inefficiency by introducing FrugalPrompt, a novel prompt compression framework for LLMs, which retains only the most semantically significant tokens. Leveraging two state-of-the-art token attribution methods, GlobEnc and DecompX, we assign salience scores to every token in an input sequence, rank them to retain the top-$k\%$ tokens, and obtain a sparse frugalized prompt. We establish the theoretical stability of our approach and provide strong empirical results across a suite of four NLP tasks to study the trade-off between the portion of retained tokens and performance. Experimental findings across retention settings reveal asymmetric performance patterns that suggest potential task contamination effects. We posit that our work contributes to a more nuanced understanding of LLM behavior in performance-efficiency trade-offs and delineates the boundary between tasks tolerant of contextual sparsity and those requiring exhaustive context.
FrugalPrompt: Reducing Contextual Overhead in Large
Language Models via Token Attribution
Syed Rifat Raiyan11, Md Farhan Ishmam22, Abdullah Al Imran33, Mohammad Ali Moni44 11Islamic University of Technology, Bangladesh  22University of Utah, USA 33University of Liverpool, UK  44University of Queensland, Australia rifatraiyan@iut-dhaka.edu, farhan.ishmam@utah.edu
1 Introduction
As large language models (LLMs) are now becoming general-purpose intelligence systems for NLP, they are supporting workflows that go beyond their initial zero-shot text generation: from coding systems capable of using tools to multi-step agentic systems solving multimodal problems Yang et al. (2024); Schmidgall et al. (2025). Yet the same empirical scaling laws that fuel these capabilities into trillion-parameter models Fedus et al. (2022); Hudson et al. (2023) also intensify a basic cognitive and computational tension. Human language is richly redundant [^1] We define redundancy as any verbiage that extends beyond the most succinct form of communication., while modern LLM systems being trained and conditioned on human text are constrained by the context length Wu et al. (2025), latency Wan et al. (2023), and energy costs Stojkovic et al. (2024) due to the expressiveness of human language.
A core observation of studying linguistics from the lens of information theory is that natural language is not information-minimal code Shannon (1951). Human utterances inherently contain around $80\%$ statistical redundancy, e.g., text is repeated for clarification, and function words support grammatically more than semantic context. Many tokens, therefore, contribute little incremental information once the local context is known. Psycholinguistic evidence similarly suggests that comprehenders can often recover intended meaning from “telegraphic” input, a property exploited in simplified registers and second-language communication Gibson et al. (2013); Klein and Perdue (1997). These findings suggest a cognitively motivated route to efficiency in NLP, and we ask: To what extent do LLMs require full linguistic explicitness, and to what extent can they reconstruct meaning from pragmatically sufficient cues?
@@ 1 Introduction
We aim to answer these questions with a theoretically grounded work motivated by information-theoretic approaches in linguistics. Using SOTA token attribution techniques, we compute saliency scores for the prompt and selectively retain the semantically rich tokens while filtering redundant ones. This yields a training-free compressor that (i) better preserves meaning-critical content, (ii) systematically removes low-value discourse material and function words, and (iii) avoids the cost of an auxiliary large model. Importantly, our approach requires only a lightweight pretrained encoder, providing substantially lower overhead than prior compressor designs. By introducing a prompt compression strategy that uses saliency scores generated from state-of-the-art pre-trained token attribution methods Modarressi et al. (2022, 2023). These saliency scores can capture the semantically rich token while filtering the redundant ones. This is achieved without any training paradigm, with minimal parameter overhead of a single 110M BERT model Devlin et al. (2019), substantially lower than that of existing methods.
@@
Our criteria for filtering are grounded in two linguistic principles underlying textual redundancy. First, there is an intrinsic redundancy to any form of text produced by humans Wit and Gillette (1999). For instance, in the text, “What are your plans for the weekend, by the way?”, the phrase “by the way” contributes no additional meaning to an LLM. Second, the English language has low-importance tokens that add little to no semantic weight, generally in the form of articles and auxiliary verbs. For instance, most of the meaning is conveyed if the statement “I have received a parcel.”, is simplified to “I received parcel.”, with the latter mimicking simplified English, often used by non-native speakers Shubert et al. (1995). Our filtering methods eliminate both forms of redundancy. Our key contributions are as follows:
@@
  • We introduce a novel, training-free prompt frugalization strategy for LLMs that controllably filters low-importance tokens based on saliency scores from pre-trained encoders.
  • Our method is theoretically grounded with empirical results showing strong retention–performance trade-offs. Across four NLP tasks, a $20\%$ prompt reduction preserves performance for most models with negligible parameter overhead.
@@
  • We reveal behavioral insights into how scaling performance relates to inference cost and potential task contamination in benchmarks.
2 Related Work
2.1 Redundancy in Language
Language, by nature, is not information-minimal code Shannon (1951) and introduces several forms of redundancy. Through the lens of information theory Shannon (1948), redundancy of a language can be defined as the difference between the maximum possible entropy of the language and its actual entropy. Theories in inferential pragmatics Richards (1990) and relevance-driven communication argue that interlocutors reconstruct implicit content by focusing on semantically informative cues rather than processing every token with equal weight. Such laconism is not an anomaly but a fundamental property of human communicative efficiency Levshina and Moran (2021).
2.2 LLM Efficiency
@@
The current literature on LLM efficiency Wan et al. (2023) stems from earlier works on efficient transformers Tay et al. (2022). Inference-time efficiency can be achieved via model compression methods, such as quantization Dettmers et al. (2023b), parameter pruning Ma et al. (2023), low-rank decomposition Li et al. (2023a), and knowledge distillation Timiryasov and Tastet (2023). Training efficiency can be achieved during pre-training Liu et al. (2023); Yao et al. (2024) and fine-tuning, either through parameter-efficient Hu et al. (2022) or memory-efficient Dettmers et al. (2023a) approaches. Finally, architectural changes also improve efficiency, e.g., efficient attention Child et al. (2019); Ainslie et al. (2023) and state-space models Gu and Dao (2023).
Distinct from these methods, our work adopts a data-centric approach, where the input prompt is optimized rather than the model’s underlying weights or architecture Chang et al. (2024). In essence, we perform prompt compression, which minimizes input redundancy while retaining task-specific information. Li et al. (2023b); Jiang et al. (2023). Contrasting previous approaches that often use a paraphrasing model Chuang et al. (2024), which can lead to instability, ours takes a theoretically grounded approach by filtering the least salient tail-end tokens.
2.3 Token Attribution
@@
The self-attention weights in transformers have been primarily explored as a proxy for token importance Htut et al. (2019); Clark (2019), yet they fail to align with the true attribution of tokens Jain and Wallace (2019). While including vector norms can partially mitigate this issue Kobayashi et al. (2020), it overlooks the intricacies of the non-attention components in a transformer. Instead, we rely on subsequent token attribution methods (see Appendix A): GlobEnc Modarressi et al. (2022), which quantifies global token attribution by aggregating layer-wise analyses across the entire encoder using attention rollout and vector norms, and DecompX Modarressi et al. (2023), which propagates locally decomposed token representations globally to incorporate all model components, including non-linear feed-forward networks. These methods were selected for their superior ability to capture comprehensive semantic contributions beyond raw attention weights, enabling identification of salient tokens with minimal computational overhead.
@@
3 Methodology
Figure 1 portrays an overview of our proposed pipeline. We obtain the saliency scores using the token attribution methods: GlobEnc Modarressi et al. (2022) and DecompX Modarressi et al. (2023), details discussed in Appendix A and employ frugalization via token filtration to generate a condensed yet semantically representative version of an input sample. This technique identifies and retains only the most salient tokens, thereby reducing input dimensionality while preserving the core meaning for the given task.
3.1 Ranking via Saliency Scores
@@ 3.2 Top-$k$ Filtering
We reindex $S_{k}$ in monotone increasing order to produce ${S}_{k\uparrow}$. The frugalized sequence ${F}_{k}$ is produced by restoring the selected tokens to their original order, i.e.,
@@
Model Token Attribu- tion $k\%$ Tasks
@@
3.3 Theoretical Bounds and Relationships
We derive the theoretical bounds of token filtering or deletion based on saliency scores. The bounds formalize the relationship between dropping $k\%$ tokens and the estimated performance.
Deletion Effect.
For a filtered or deleted set of indices $D\subseteq[n]=\{1,\dots,n\}$, we define the retained subsequence of tokens $T\setminus D:=T_{[n]\setminus D}$. For task $\tau$ and model parameters $\theta$, let $\ell_{\tau}(T,y;\theta)\in\mathbb{R}^{+}$ be the task loss or evaluation score. We define the deletion-effect set function:
$f(D)\;:=\;\underbrace{\ell_{\tau}(T\setminus D,y;\theta)}_{\text{Loss after Deletion}}-\underbrace{\ell_{\tau}(T,y;\theta)}_{\text{Original Loss}}.$ (1)
Trivially, $f(\emptyset)=0$. Eq. 1 is intuitively straightforward; $f(D)>0$ means deletion worsens the performance, whereas $f(D)<0$ means deletion improves performance, which is plausible. We emphasize that all results are inference-time, i.e., the model parameters $\theta$ are fixed. We use $\ell_{T}(T,y;\theta)$ solely as an evaluation functional to quantify how token deletion affects predictive performance.
Discrete Marginals and Interactions.
For a set of deleted tokens $A=[n]=\subseteq\{1,\dots,n\}$ and a new token $i\notin A$, we define marginal deletion:
$\Delta_{i}f(A)\;:=\;f(A\cup\{i\})-f(A).$ (2)
For tokens $i\neq j$ with $i,j\notin A$, we define the second-order discrete difference, i.e., the pairwise interaction as
$\Delta_{ij}f(A)\;:=\;\Delta_{i}\Delta_{j}f(A)=f(A\cup\{i,j\})$
$-f(A\cup\{i\})-f(A\cup\{j\})+f(A).$ (3)
Assumption 1 (Saliency upper-bounds singleton deletion harm).
There exists a task constant, $C_{\tau}\geq 0$, such that,
$f(\{i\})\;$ $\leq\;C_{\tau}\,s_{i},\quad\forall i\in[n].$ (4)
Assumption 1 is essentially true, as deleting token $i$ alone cannot increase the deletion effect more than a constant factor of its saliency.
Assumption 2 (Bounded pairwise interaction).
There exists a task constant, $\gamma_{\tau}\geq 0$, such that,
$\Delta_{ij}f(A)\;$ $\leq\;\gamma_{\tau},\quad\forall A\subseteq[n],\forall i,j\notin A.$ (5)
Assumption 2 simply imposes a pairwise bound that prevents the deletion effect of multiple tokens from causing the harm to explode.
Theorem 1 (Deletion bound via saliency and interaction).
Under Assumption 1, 2, for any deletion set $D$ with $|D|=q$,
$f(D)\;\leq\;C_{\tau}\sum_{i\in D}s_{i}\;+\;\frac{\gamma_{\tau}}{2}\,q(q-1).$ (6)
The proof of Theorem 1 is provided in Appendix §B. Theorem 1 establishes the stability of token deletion by upper-bounding the deletion effect.
Retention Relationship.
Following §3.2, we define the deleted set $D_{k}:=\{\pi(p+1),\dots,\pi(n)\}$ and the frugalized sequence $F_{k}:=T\setminus D_{k}$.
Corollary 1 (Performance vs. $k\%$ retention).
Under Assumption 1, with q=npq=n-p,
$\ell_{\tau}(F_{k},y;\theta)-\ell_{\tau}$ $(T,y;\theta)\;\leq\;C_{\tau}\sum_{j=p+1}^{n}s_{\pi(j)}\;$
$+\;\frac{\gamma_{\tau}}{2}\,(n-p)(n-p-1).$ (7)
Corollary 1 can be derived from Eq. 1, 5, and can also be generalized at the dataset level by computing the risk. Given $(T,y)\sim\mathcal{D}_{\tau}$ denotes the test distribution for task $\tau$, the risk can be simply defined as the expectation over the loss:
$R_{\tau}(\theta):=\mathbb{E}_{(x,y)\sim\mathcal{D}_{\tau}}[\ell_{\tau}(x,y;\theta)]$ (8)
Figure 2: The performance difference between the reduced tokens and baseline performance using $100\%$ tokens of metrics across three tasks: text classification (CLS), question-answering (QA), and reasoning (RSN).
4 Experiment
With the performance bounds and relationship with the retention kk% are strongly conveyed in §3.3, we provide empirical evidence in this section.
4.1 Tasks
@@
4.2 Dataset
For sentiment analysis, we use IMDb movie review Maas et al. (2011), a binary sentiment analysis corpus comprising movie reviews labeled positive or negative. We use Argilla News to evaluate abstractive summarization on news articles Ahmed et al. (2017). For commonsense QA, we use CosmosQA Huang et al. (2019), which primarily focuses on commonsensical and causal reasoning within short narrative passages. The questions stipulated in CosmosQA are easy for humans to solve, but usually challenging for LLMs. We use GSM8k Cobbe et al. (2021) for mathematical reasoning, which contains grade-school MWPs to assess basic arithmetic reasoning and multi-step logical deduction.
Figure 3: The performance difference between the reduced tokens and baseline performance using $100\%$ tokens across the summarization metrics.
4.3 Models
We perform inference using a set of contemporary frontier models, varying in architecture and scale, to examine the generalizability of our method. The models are: LLaMA-3 8B, LLaMA-3 70B Meta (2024), GPT-3.5 Brown et al. (2020), Gemini 2.0 Flash Thinking Google DeepMind (2025), and o3-mini OpenAI (2025). The closed-source models are accessed via their respective APIs, and the open-source models are accessed via OpenRouter [^2] https://openrouter.ai/, using standardized prompts to minimize formatting-induced variability.
4.4 Experimental Setup
We set the temperature to 0 for all inference runs to sample from the model’s unaltered probability distribution, ensuring a standard and reproducible generation process. For the frugalized setup, we experiment by filtering the top $k\in\{50,60,80\}\%$ of the input tokens. We also include the baseline prompt, using $100\%$ of the tokens to set a proper frame of reference for the performance delta across all models. This setup enables us to investigate whether low-attribution filler tokens can be excluded with minimal degradation in task performance, thereby reducing API costs and carbon footprint.
4.5 Evaluation Metrics